OpenStudy (anonymous):

Select one of the factors of x2 + 7x – 44 Answer x + 2 x – 11 x – 4 x – 22 .

jhonyy9 (jhonyy9):

(x+11)(x-4)=x*2 +11x-4x-44=x*2 +7x -44 - so the answer x-4